Find the unknown emf and the unknown resistor in t
Answer #1

Apply node equation at point C.

Current through resistor R2 will be (I1+I2).

Apply loop equation to the left loop -

125 - E = R2(I1+I2) + I1*R1 = 3(2.5+9.0) + 2.5*7.0 = 52

=> E = 125-52 = 73 V

Now, apply loop equation in the right loop -

125 = R2(I1+I2) - I2*R = 3(2.5+9) + 9*R

=> 9R = 125 - 34.5 = 90.5

=> R = 10.05 Ohm